# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix D Proof of Lipschitz-continuity Bound
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)-\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}(x)\bigg{|}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\right|$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}\right|$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right|$
The second equality uses Lemma D.8. The third equality uses the property that each local agent can only have one state/action pair, and the fourth equality vectorizes the indicator random variables. Then,
@@
$\quad\quad\quad\quad\leq\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right|$
@@
$\quad\quad\quad\quad=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}$
@@
$\quad\quad\quad\quad\leq\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}\cdot\|\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\|_{1}$
@@
$\quad\quad\quad\quad=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right\|_{1}$
@@
$\quad\quad\quad\quad=\frac{1}{2}\sum_{(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right|$
$\quad\quad\quad\quad=\frac{1}{2}\sum_{z_{l}\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|F_{z_{\Delta}}(z_{l})-F_{z_{\Delta^{\prime}}}(z_{l})\right|$
$\quad\quad\quad\quad:=\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
@@
The first inequality uses the triangle inequality. The second inequality and fifth equality follow from Hölder’s inequality and the sum of the probabilities from the stochastic transition function $\mathcal{J}_{1}$ being equal to $1$. The sixth equality uses Fubini-Tonelli’s theorem which applies as the total variation distance measure is bounded from above by $1$. The final equality recovers the total variation distance for the variable $z=(s,a)\in\mathcal{S}_{l}\times\mathcal{A}_{l}$ across agents $\Delta$ and $\Delta^{\prime}$.∎
Lemma D.8.
@@
$\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)$ $=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$=\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})$
The second equality follows from the conditional independence of $s_{i}^{\prime}$ from $s_{\Delta\cup\Delta^{\prime}\setminus i}$ from Lemma D.9, and the final equality uses the fact that the expectation of an indicator random variable is the probability distribution function of the random variable. ∎
Lemma D.9.
@@
We direct the interested reader to the Bayes-Ball theorem in Shachter (2013) for proving conditional independence. For ease of exposition, we restate the two rules in Shachter (2013) that introduce the notion of $d$-separations which implies conditional independence. Suppose we have a causal graph $G=(V,E)$ where the vertex set $V=[p]$ for $p\in\mathbb{N}$ is a set of variables and the edge set $E\subseteq 2^{V}$ denotes dependence through connectivity. Then, the two rules that establish $d$-separations are as follows:
@@
Therefore, $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are $d$-separated by $Z$. Hence, by Shachter (2013), $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are conditionally independent.∎
Lemma D.10.
@@
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$\quad\quad\quad=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
@@
We start by expanding the expectations:
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$\quad\quad\quad=\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]\max_{\begin{subarray}{c}a_{g}^{\prime\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$\quad\quad\quad=\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}^{2}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}^{T}_{k}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$\quad\quad\quad=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
In the second equality, we note that the right-stochasticity of $\mathcal{J}_{k}$ implies the right-stochasticity of $\mathcal{J}_{k}^{2}$. Further, observe that $\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]$ denotes the probability of the transitions $(s_{g},s_{\Delta})\to(s_{g}^{\prime},s_{\Delta}^{\prime})\to(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ with actions $a_{g},a_{\Delta}$ at each step, where the joint state evolution is governed by $\mathcal{J}_{k}$. Thus,

### v3

@@ Appendix E Proof of Lipschitz-continuity Bound
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}(x)\bigg{|}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\bigg{|}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}\bigg{|}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\bigg{|}$
The second equality uses Lemma E.9. The third equality uses the property that each local agent can only have one state/action pair, and the fourth equality vectorizes the indicator variables. Then,
@@
$\leq\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right|$
@@
$=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}$
@@
$\leq\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}\cdot\|\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\|_{1}$
@@
$=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right\|_{1}$
@@
$=\frac{1}{2}\sum_{(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right|$
$=\frac{1}{2}\sum_{z_{l}\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|F_{z_{\Delta}}(z_{l})-F_{z_{\Delta^{\prime}}}(z_{l})\right|$
$:=\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
@@
The first inequality uses the triangle inequality. The second inequality and fifth equality follow from Hölder’s inequality and the sum of the probabilities from the stochastic transition function $\mathcal{J}_{1}$ being equal to $1$. The sixth equality uses Fubini-Tonelli’s theorem which applies as the total variation distance measure is bounded from above by $1$. The final equality recovers the total variation distance for the variable $z=(s,a)\in\mathcal{S}_{l}\times\mathcal{A}_{l}$ across agents $\Delta$ and $\Delta^{\prime}$, which proves the claim.∎
Next, recall the data processing inequality.
Lemma E.8 (Data Processing Inequality.).
Let $A$ and $B$ be random variables over some domain $S$. Let $f$ be some function (not necessarily deterministically) mapping from $S$ to any codomain $T$. Then, every $f$-divergence $\chi$ satisfies
$\chi(f(A)\|f(B))\leq\chi(A\|B)$
Remark: Note that TV-distance is not an $f$-divergence, and therefore it does not generally satisfy the data processing inequality. Under this lens, Lemma E.7 is novel since it saturates an analogous data-processing relation for the TV distance in our multi-agent setting.
Lemma E.9.
@@
$\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)$
$\quad\quad\quad=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$\quad\quad\quad=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$\quad\quad\quad=\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})$
The second equality follows from the conditional independence of $s_{i}^{\prime}$ from $s_{\Delta\cup\Delta^{\prime}\setminus i}$ from Lemma E.10, and the final equality uses the fact that the expectation of an indicator random variable is the probability distribution function of the random variable. ∎
Lemma E.10.
@@
We direct the interested reader to the Bayes-Ball theorem in Shachter [2013] for proving conditional independence. For ease of exposition, we restate the two rules in Shachter [2013] that introduce the notion of $d$-separations which implies conditional independence. Suppose we have a causal graph $G=(V,E)$ where the vertex set $V=[p]$ for $p\in\mathbb{N}$ is a set of variables and the edge set $E\subseteq 2^{V}$ denotes dependence through connectivity. Then, the two rules that establish $d$-separations are as follows:
@@
Therefore, $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are $d$-separated by $Z$. Hence, by Shachter [2013], $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are conditionally independent.∎
Lemma E.11.
@@
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$\quad\quad\quad=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g}}\hat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
@@
By expanding the expectations:
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$=\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\max_{\begin{subarray}{c}a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$=\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}^{2}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}^{T}_{k}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\hat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
In the second equality, the right-stochasticity of $\mathcal{J}_{k}$ implies the right-stochasticity of $\mathcal{J}_{k}^{2}$.
Further, observe that $\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]$ denotes the probability of the transitions $(s_{g},s_{\Delta})\to(s_{g}^{\prime},s_{\Delta}^{\prime})\to(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ with actions $a_{g},a_{\Delta}$ at each step, where the joint state evolution is governed by $\mathcal{J}_{k}$. Thus,